# Deep Reinforcement Learning for Traveling Purchaser Problems

### v5

@@ Abstract
The traveling purchaser problem (TPP) is an important combinatorial optimization problem with broad applications. Due to the coupling between routing and purchasing, existing works on TPPs commonly address route construction and purchase planning simultaneously, which, however, leads to exact methods with high computational cost and heuristics with sophisticated design but limited performance. In sharp contrast, we propose a novel approach based on deep reinforcement learning (DRL), which addresses route construction and purchase planning separately, while evaluating and optimizing the solution from a global perspective. The key components of our approach include a bipartite graph representation for TPPs to capture the market-product relations, and a policy network that extracts information from the bipartite graph and uses it to sequentially construct the route. One significant benefit of our framework is that we can efficiently construct the route using the policy network, and once the route is determined, the associated purchasing plan can be easily derived through linear programming, while, leveraging DRL, we can train the policy network to optimize the global solution objective. Furthermore, by introducing a meta-learning strategy, the policy network can be trained stably on large-sized TPP instances, and generalize well across instances of varying sizes and distributions, even to much larger instances that are never seen during training. Experiments on various synthetic TPP instances and the TPPLIB benchmark demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics, reducing the optimality gap by 40%-90%, and also showing an advantage in runtime, especially on large-sized instances.
@@ 1 Introduction
Unfortunately, TPPs are known to be strongly NP-hard (Laporte et al., 2003). In particular, the main challenge in solving TPPs stems from the inherent coupling between routing and purchasing: the routing decisions have to take both traveling and purchasing costs into consideration, while the choice of visited markets will, in turn, impact the purchasing decisions and associated cost. Thus, existing works typically need to deal with both routing and purchasing simultaneously (Manerba et al., 2017), which, however, imposes many limitations. On the one hand, exact methods such as branch-and-cut (Laporte et al., 2003; Riera-Ledesma and González, 2006) require joint optimization of routing and purchasing decisions, which generically leads to high computational cost, making them often intractable for realistically sized problem instances. For example, solving a medium-sized TPP instance with 100 markets and 50 products can take hours on a computer with a 2.3 GHz processor, and the computational time grows exponentially with the problem size (Riera-Ledesma, 2012). On the other hand, various heuristic methods, such as the Generalized Savings Heuristic (GSH) (Golden et al., 1981), the Tour Reduction Heuristic (TRH) (Ong, 1982), and the Commodity Adding Heuristic (CAH) (Pearn and Chien, 1998), have been proposed to produce high-quality solutions for TPPs within a reasonable time. Readers are referred to Manerba et al. (2017) for further details. However, these heuristics have to carefully balance the effect of each operation on the traveling cost and purchasing cost, and thus rely heavily on sophisticated design that requires substantial human expertise and domain knowledge. Moreover, these hand-crafted heuristics typically need to be tailored case-by-case, resulting in that they may only be effective on limited instances with specific characteristics, and lack the ability to generalize across different instance distributions. For example, CAH performs poorly if the purchasing cost dominates in the optimization objective, and GSH may yield poor solutions if a market sells most of the products but is located far from other markets (Pearn and Chien, 1998).
@@
In this paper, we propose a novel approach based on deep reinforcement learning (DRL), which addresses the limitations of existing methods by decoupling the treatment of routing and purchasing decisions. The core idea of our approach can be summarized as “solve separately, learn globally”. We note that by leveraging the forward-thinking and global-optimizing mechanisms of DRL, we can break the complex task of solving TPPs into two separate stages: route construction and purchase planning, while learning a policy to guide the decision-making towards optimizing the global solution objective. Specifically, in the first stage, we use a policy network to sequentially construct the route, where at each decision step the policy network takes the problem instance and current partial route as input, and outputs an action distribution to determine the next market to visit. Then, once a complete route is constructed, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets through an easily solvable linear transportation problem. Note that in the first stage, we use the policy network only to decide the route, i.e., which markets to visit and the visited order, deferring the decisions on specific purchasing plan to the second stage. This separation decouples the routing decisions and purchasing decisions at the operating level, such that each stage can be efficiently done during the solving process. Meanwhile, to bridge the optimization interdependence between routing and purchasing, we train the first-stage policy network to optimize the global solution objective, which is determined jointly by both stages. In other words, though the policy network is only used for route construction in the first stage, it is learned to construct a high-quality route that can not only have low traveling cost, but also be promising to lead to a low-cost purchasing plan in the subsequent second stage, thus to minimize the overall solution objective.
@@
In fact, the idea of leveraging DRL for CO problems is not entirely new though (Lopes Silva et al., 2019; Bengio et al., 2021; Que et al., 2023; Fajemisin et al., 2024), particularly in routing problems such as the traveling salesman problem (TSP) (Vinyals et al., 2015; Bello et al., 2017; Zhang et al., 2023; Gao et al., 2024) and the vehicle routing problem (VRP) (Nazari et al., 2018; Zhang et al., 2020; Kalatzantonakis et al., 2023; Rodríguez-Esparza et al., 2024). However, these efforts typically rely on problem-specific properties and are limited to problems with simple structures, which cannot be readily extended to TPPs. Therefore, considering the double nature of TPPs and building on our solution framework, we first propose a novel bipartite graph representation for TPPs, where the markets and products are represented as two sets of nodes, with the supply information (e.g., the supply quantities and prices) encoded as edge features between the market and product nodes. In contrast to existing representations for routing problems (e.g., the complete or k-NN graphs (Kool et al., 2019; Fu et al., 2021)), our bipartite graph representation can naturally capture the relations between markets and products through message-passing along edges. Then, we design a policy network, which can effectively extract information from the bipartite graph and use it for route construction. Specifically, our policy network exploits the connecting structure of the bipartite graph, and leverages graph neural networks (GNNs) and multi-head attention (MHA) to aggregate information both across and within the two sets of nodes. This effectively facilitates the extraction of relational information of the markets and products, such as the spatial relations between markets and the potential substitutions and complementarities in product supply, which is important for constructing a high-quality route. Moreover, it is worth mentioning that our bipartite graph representation and policy network have another appealing property: they are size-agnostic such that a trained policy network can be flexibly adapted to TPP instances of varying sizes without the need for retraining or parameter adjustment.
@@
In addition, despite our proposed framework and the key DRL components introduced above, how to efficiently learn a high-quality policy remains another challenge. Due to the huge state-action space, a randomly initialized policy network may suffer from inefficient exploration, which can result in slow convergence or even training collapse, particularly on large-sized instances. Different from the training techniques or tricks (Kwon et al., 2020; Kim et al., 2022) designed for specific problems (mainly TSP or VRP), we propose an effective and more generalizable training strategy based on meta-learning (Finn et al., 2017). The strategy trains an initialized policy network on a collection of small-sized instances, with the learning objective to achieve fast adaptation to large-sized instances at low fine-tuning cost. Furthermore, we observe that the initialized policy network can learn cross-distribution knowledge through meta-learning, such that it can effectively generalize across varied instance sizes and distributions, even achieving good zero-shot generalization performance to instances of much larger sizes that are never seen during training.
@@ 2 Problem Formulation
Mathematically, a TPP is defined on a complete directed graph $\mathcal{G}=\left(V,E\right)$, where $V=\left\{v_{0}\right\}\cup M$ is the set of nodes and $E=\{(i,j):i,j\in V,i\neq j\}$ is the set of edges. Node $v_{0}$ denotes the depot and $M$ denotes the set of markets. The problem considers a set $K$ of products to purchase, where a demand $d_{k}$ is specified for each product $k\in K$. Each product $k$ is supplied in a subset of markets $M_{k}\subseteq M$, where at most $q_{ik}$ units of product $k$ can be purchased from market $i\in M_{k}$, at a price of $p_{ik}$. The traveling cost associated with each edge $(i,j)\in E$ is denoted by $c_{ij}$. The purchaser needs to determine a route on $\mathcal{G}$, i.e., a simple cycle $\pi=\left(v_{0},v^{1},v^{2},\dots,v^{T},v_{0}\right),\;T\leq|M|$, which starts and ends at the depot and visits a subset of markets, and decide how much of each product to purchase at each market to satisfy the demands at minimum traveling and purchasing costs. Note that to guarantee the existence of a feasible purchasing plan with respect to the product demand, it is assumed that $0<q_{ik}\leq d_{k}$ for each $k\in K$ and $i\in M_{k}$, and $\sum_{i\in M_{k}}q_{ik}\geq d_{k}$ for each $k\in K$. In the case of U-TPP, the first assumption becomes $q_{ik}=d_{k}$ for all $k\in K$ and $i\in M_{k}$, i.e., the supply quantities for the available products at each market are unlimited, and otherwise R-TPP.
@@
$\delta^{-}\left(V^{\prime}\right):=\left\{(i,j)\in E:i\notin V^{\prime},j\in V^{\prime}\right\}.$
Then, the TPP can be formulated as follows (Laporte et al., 2003):
@@
The MILP formulation (1)-(9) is commonly employed in exact methods for TPPs (Manerba et al., 2017). However, the number of constraints is exponential in the number of markets and the number of variables is also significantly large, which can lead to a massive branch-and-bound tree with weak lower bounds in an MILP solver. The branch-and-cut algorithms, as introduced in Laporte et al. (2003) and Riera-Ledesma and González (2006), attempt to alleviate this issue by dynamically generating variables and separating constraints, thereby to reduce the tree size and accelerate solving. Nonetheless, the dynamic pricing and separation procedures are still computationally expensive, making it often unaffordable to solve the MILP exactly in practice, especially for large-sized TPP instances. In contrast, our DRL-based approach does not directly solve the MILP, but instead creates a bipartite graph representation (Section 4) based on its structure, which is further used as input to the policy network for route construction.
@@
We now introduce our “solve separately, learn globally” framework for TPPs. As aforementioned, prior works typically require to address routing and purchasing simultaneously (Manerba et al., 2017). In contrast, by leveraging DRL, we can decouple the routing decisions and purchasing decisions at the operating level, i.e., “solve separately”, while guiding the decision-making using a policy, which is learned to optimize the global solution objective, i.e., “learn globally”. This framework offers the potential for both high computational efficiency and high-quality solutions if the policy is well-trained.
Figure 1: Our “solve separately, learn globally” framework. In the first stage, the purchaser starts from the depot, and selects a new node to add to the partial route based on a policy network at each decision step, until it returns to the depot and a complete route is made. Then in the second stage, the purchasing plan is derived through a transportation problem, which is integrated into the reward calculation procedure for the terminal state. The final reward is defined as the negative solution objective, which is used to update the policy network through a DRL training algorithm.
@@ 3.1 Overview
Figure 1 illustrates an overview of our proposed framework. The solving process consists of two separate stages: route construction and purchase planning. In implementation, we formulate this process as a Markov decision process (MDP), which provides a fundamental framework for DRL. Specifically, in the first stage, we sequentially construct the route through finite decision steps. The purchaser starts from the depot as the initial state, and at each decision step selects the next market (or the depot) to visit based on a policy, which is parameterized as a deep neural network (called the policy network). The route construction process iterates until the depot is revisited, i.e., a complete route is formed, and the MDP reaches the terminal state. Then, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets along the route by solving a linear transportation problem. In our implementation of MDP, the second stage is integrated into the reward calculation procedure for the terminal state. The final reward is defined as the negative solution objective, which is the sum of the traveling cost of the route constructed in the first stage and the purchasing cost derived in the second stage. The reward is then used to update the first-stage policy network through a DRL training algorithm.
@@ 3.2.1 State
In the MDP, a state should contain necessary information that the policy network needs for decision-making at the current step. In our task for route construction, the state at decision step $t$, denoted as $s_{t}$, consists of two parts: 1) the TPP instance $U$ being solved, which is static throughout the solving process, and 2) the dynamic information of the partial route that has been constructed up to step $t$. Specifically, the TPP instance $U$ is represented as a bipartite graph, which will be introduced in detail in Section 4. For the dynamic information, we consider two contents that are critical to the decision task at step $t$:
@@
    the nodes on the partial route $\pi^{t}=\left(v_{0},v^{1},v^{2},\dots,v^{t-1}\right)$,
@@
We remark that ideally, a well-trained policy network can automatically extract information about the remaining demand $d^{t}$ given the instance $U$ and the current partial route $\pi^{t}$, but we found that explicitly providing $d^{t}$ to the policy network can effectively reduce the computational load and accelerate training. Therefore, we still include the remaining demand $d^{t}$ as part of a state $s_{t}$.
@@ 3.2.2 Action
At each decision step $t$, an action $a_{t}\in A_{t}$ is defined as selecting the next node (i.e., a new market or the depot) to visit. If a new market is selected, the purchaser will visit it and start the next step $t+1$. Otherwise, if the depot is selected to visit again, a complete route is formed and the route construction terminates.
Note that an arbitrarily selected action may lead to infeasible solutions, either because 1) the constructed route may not be a simple cycle, or 2) the markets on the route cannot provide any feasible purchasing plan that fulfills the product demands in the second stage. Therefore, to avoid infeasibility, we introduce two masking rules over the action set $A_{t}$ at each decision step $t$. First, in order to strictly form a simple cycle, markets already in the partial route $\pi^{t}$ are excluded from the action set $A_{t}$. In fact, it makes no sense to revisit a market from the perspective of optimization objective: the traveling cost will increase strictly, but no savings on the purchasing cost can be made by revisiting a market. Second, to guarantee the existence of a feasible purchasing plan based on the route, we force the depot to be masked from $A_{t}$ until the remaining demand $d_{k}^{t}=0$ for all $k\in K$. In other words, the purchaser is forbidden to return to the depot unless all product demands can possibly be satisfied from the markets on the current partial route. With these two masking rules, our approach is guaranteed to produce a feasible route and purchasing plan regardless of the quality of policy.

### v6

@@ Abstract
The traveling purchaser problem (TPP) is an important combinatorial optimization problem with broad applications. Due to the coupling between routing and purchasing, existing works on TPPs commonly address route construction and purchase planning simultaneously, which, however, leads to exact methods with high computational cost and heuristics with sophisticated design but limited performance. In sharp contrast, we propose a novel approach based on deep reinforcement learning (DRL), which addresses route construction and purchase planning separately, while evaluating and optimizing the solution from a global perspective. The key components of our approach include a bipartite graph representation for TPPs to capture the market-product relations, and a policy network that extracts information from the bipartite graph and uses it to sequentially construct the route. One significant advantage of our framework is that we can efficiently construct the route using the policy network, and once the route is determined, the associated purchasing plan can be easily derived through linear programming, while, by leveraging DRL, we can train the policy network towards optimizing the global solution objective. Furthermore, by introducing a meta-learning strategy, the policy network can be trained stably on large-sized TPP instances, and generalize well across instances of varying sizes and distributions, even to much larger instances that are never seen during training. Experiments on various synthetic TPP instances and the TPPLIB benchmark demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics, reducing the optimality gap by 40%-90%, and also showing an advantage in runtime, especially on large-sized instances.
@@ 1 Introduction
Unfortunately, TPPs are known to be strongly NP-hard (Manerba et al., 2017). In particular, the main challenge in solving TPPs stems from the inherent coupling between routing and purchasing: the routing decisions have to take both traveling and purchasing costs into consideration, while the choice of visited markets will, in turn, impact the purchasing decisions and associated costs. Thus, existing works typically need to deal with both routing and purchasing simultaneously (Manerba et al., 2017), which, however, imposes many limitations. On the one hand, exact methods such as branch-and-cut (Laporte et al., 2003; Riera-Ledesma and González, 2006) require joint optimization of routing and purchasing decisions, which generically leads to high computational cost, making them often intractable for realistically sized problem instances. For example, solving a medium-sized TPP instance with 100 markets and 50 products can take hours on a computer with a 2.3 GHz processor, and the computational time grows exponentially with the problem size (Riera-Ledesma, 2012). On the other hand, various heuristic methods, such as the Generalized Savings Heuristic (GSH) (Golden et al., 1981), the Tour Reduction Heuristic (TRH) (Ong, 1982), and the Commodity Adding Heuristic (CAH) (Pearn and Chien, 1998), have been proposed to produce high-quality solutions within a reasonable time. Readers are referred to the Appendices for further details. However, these heuristics have to carefully balance the effect on traveling and purchasing costs for each operation, and thus rely heavily on sophisticated design that requires substantial human expertise and domain knowledge. Moreover, these hand-crafted heuristics typically need to be tailored case-by-case, resulting in that they may only be effective on limited instances with specific characteristics, and lack the ability to generalize across different instance distributions. For example, CAH performs poorly if the purchasing cost dominates in the optimization objective, and GSH may yield poor solutions if a market sells most products but is located far from other markets (Pearn and Chien, 1998).
@@
In this paper, we propose a novel approach based on deep reinforcement learning (DRL), which addresses the limitations of existing methods by decoupling the treatment of routing and purchasing decisions. The core idea of our approach can be summarized as “solve separately, learn globally”. We notice that by leveraging the forward-thinking and global-optimizing mechanisms of DRL, we can break the complex task of solving TPPs into two separate stages: route construction and purchase planning, while learning a policy to guide the decision-making towards optimizing the global solution objective. Specifically, in the first stage, we use a policy network to sequentially construct the route, where at each decision step the policy network takes the problem instance and current partial route as input, and outputs an action distribution to determine the next market to visit. Then, once a complete route is constructed, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets through an easily solvable linear transportation problem. Note that in the first stage, we use the policy network only to decide the route, i.e., which markets to visit and the visited order, deferring the decisions on specific purchasing plan to the second stage. This separation decouples the routing decisions and purchasing decisions at the operational level, such that each stage of decisions can be efficiently done. Meanwhile, to bridge the optimization interdependence between routing and purchasing, we train the first-stage policy network to optimize the global solution objective, which is determined jointly by both stages. In other words, though the policy network is only used for route construction in the first stage, it is learned to construct a high-quality route that can not only have low traveling cost, but also be promising to lead to a low-cost purchasing plan in the second stage, thus to minimize the global solution objective.
@@
In fact, the idea of leveraging DRL for CO problems is not entirely new though (Bengio et al., 2021; Tian et al., 2023; Xin et al., 2021; Luo et al., 2024), particularly in routing problems such as the traveling salesman problem (TSP) (Vinyals et al., 2015; Bello et al., 2017; Zhang et al., 2023; Gao et al., 2024) and the vehicle routing problem (VRP) (Nazari et al., 2018; Zhang et al., 2020; Yuan et al., 2023; Wang et al., 2025). However, these efforts typically rely on problem-specific properties and are limited to problems with simple structures, which cannot be readily extended to TPPs. Therefore, considering the double nature of TPPs, we first propose a novel bipartite graph representation as an input to the policy network, where the markets and products are represented as two sets of nodes, with the supply information (e.g., the supply quantities and prices) encoded as edge features between the market and product nodes. In contrast to existing representations for routing problems (e.g., the complete or k-NN graphs (Kool et al., 2019; Joshi et al., 2019)), our bipartite graph representation can naturally capture the relations between markets and products through message-passing along edges. Then, we design a policy network, with an architecture that can effectively extract information from the bipartite graph and use it for route construction. Specifically, our policy network exploits the connecting structure of the bipartite graph, and leverages graph neural networks (GNNs) and multi-head attention (MHA) to aggregate information both within and across the two sets of nodes. This effectively facilitates the extraction of relational information of the markets and products, such as the spatial relations between markets and the potential substitutions and complementarities in product supply, which is important for the construction of high-quality routes. Moreover, it is worth mentioning that our bipartite graph representation and policy network are size-agnostic, such that a trained policy network can be flexibly adapted to TPP instances of varying sizes without the need for retraining or parameter adjustment.
@@
In addition, despite our proposed framework and the key components introduced above, how to efficiently learn an effective policy remains another challenge. Due to the huge state-action space, a randomly initialized policy network may suffer from inefficient exploration, which can result in slow convergence or even training collapse, particularly on large-sized instances. Different from the training techniques or tricks designed for specific problems (mainly TSP or VRP) (Kwon et al., 2020; Kim et al., 2022) , we introduce an effective and general training strategy based on meta-learning (Finn et al., 2017a), which trains an initialized policy network on a collection of varying instance distributions, with the learning objective to achieve efficient adaptation to new instances at low fine-tuning cost. Furthermore, the initialized policy network can learn cross-distribution knowledge through meta-learning, such that it can effectively generalize across varied instance sizes and distributions, even demonstrating zero-shot generalization ability to much larger instances that are never seen during training.
@@ 2 Problem Formulation
Mathematically, a TPP is defined on a complete directed graph $\mathcal{G}=\left(V,E\right)$, where $V=\left\{v_{0}\right\}\cup M$ is the set of nodes and $E=\{(i,j):i,j\in V,i\neq j\}$ is the set of edges. Node $v_{0}$ denotes the depot and $M$ denotes the set of markets. The problem considers a set $K$ of products to purchase, where a demand $d_{k}$ is specified for each product $k\in K$. Each product $k$ is supplied in a subset of markets $M_{k}\subseteq M$, where at most $q_{ik}$ units of product $k$ can be purchased from market $i\in M_{k}$, at a price of $p_{ik}$. The traveling cost associated with each edge $(i,j)\in E$ is denoted by $c_{ij}$. The goal is to determine a route on $\mathcal{G}$, i.e., a simple cycle $\pi=\left(v_{0},v^{1},v^{2},\dots,v^{T},v_{0}\right),\;T\leq|M|$, which starts and ends at the depot and visits a subset of markets, and decide how much of each product to purchase at each market to satisfy the demands at minimum traveling and purchasing costs. Note that to guarantee the existence of a feasible purchasing plan, it is assumed that $0<q_{ik}\leq d_{k}$ for each $k\in K$ and $i\in M_{k}$, and $\sum_{i\in M_{k}}q_{ik}\geq d_{k}$ for each $k\in K$. In the case of U-TPP, the first assumption becomes $q_{ik}=d_{k}$ for all $k\in K$ and $i\in M_{k}$, i.e., the supply quantities for the available products at each market are unlimited. Otherwise, the problem is referred to as R-TPP.
@@
$\delta^{-}\left(V^{\prime}\right):=\left\{(i,j)\in E:i\notin V^{\prime},j\in V^{\prime}\right\},$
where $\delta^{+}\left(V^{\prime}\right)$ and $\delta^{-}\left(V^{\prime}\right)$ denote the edges going out of and into the given node set $V^{\prime}$, respectively. Then, the TPP can be formulated as follows (Laporte et al., 2003):
@@
The MILP formulation (1)-(9) is commonly employed in exact methods for TPPs (Manerba et al., 2017). However, the number of constraints is exponential in the number of markets, and the number of variables is also significantly large, which can lead to a massive branch-and-bound tree with weak lower bounds in an MILP solver. The branch-and-cut algorithms, as introduced in Laporte et al. (2003) and Riera-Ledesma and González (2006), attempt to alleviate this by dynamically generating variables and separating constraints, thereby to reduce the tree size and accelerate solving. Nonetheless, the dynamic pricing and separation procedures are still computationally expensive, making it often unaffordable to exactly solve the problem in practice, especially for large-sized TPP instances. In contrast, our DRL-based approach does not directly solve the MILP, but instead creates a bipartite graph representation based on its structure (Section 4), which is further used as input to the policy network for route construction.
@@
We introduce our “solve separately, learn globally” framework in this section. As aforementioned, prior works typically require addressing routing and purchasing simultaneously (Manerba et al., 2017). In contrast, by leveraging DRL, we can decouple the routing decisions and purchasing decisions at the operational level, i.e., “solve separately”, while guiding the decision-making using a policy that is learned to optimize the global solution objective, i.e., “learn globally”. This framework offers the potential for both high computational efficiency and solution quality, provided the policy is well-trained.
Figure 1: Our “solve separately, learn globally” framework. In the first stage, the purchaser starts from the depot, and selects a new node to add to the partial route at each decision step until it returns to the depot. Then in the second stage, the purchasing plan is derived using a transportation problem, which is integrated into the reward calculation procedure for the terminal state. The final reward, defined as the negative solution objective value, is used to update the policy network through a DRL training algorithm.
@@ 3.1 Overview
Figure 1 illustrates an overview of our proposed framework. The solving procedure consists of two separate stages: route construction and purchase planning. In implementation, we formulate this process as a Markov decision process (MDP), based on which we use DRL to guide the decisions. Specifically, in the first stage, we sequentially construct the route through finite decision steps. The purchaser starts from the depot as the initial state, and at each decision step selects the next market (or the depot) to visit based on a policy, which is parameterized as a deep neural network (called the policy network). The route construction process iterates until the depot is revisited, indicating a complete route is formed, and the MDP reaches the terminal state. Then, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets by solving a linear transportation problem. Within the MDP, the second stage is integrated into the reward calculation procedure for the terminal state. The final reward is defined as the negative solution objective value, i.e., the sum of the traveling cost of the route constructed in the first stage and the purchasing cost derived in the second stage. The reward is then used to update the first-stage policy network through a DRL training algorithm.
@@ 3.2.1 State
A state should contain necessary information that the policy network needs for decision-making at the current step. In our MDP, the state at decision step $t$, denoted as $s_{t}$, consists of two parts: 1) the TPP instance $U$ being solved, which is static throughout the solving process, and 2) the dynamic information of the partial route that has been constructed up to step $t$. Specifically, the TPP instance $U$ is represented as a bipartite graph, which will be introduced in detail in Section 4. For the dynamic information, we consider two contents that are critical to the decision task at step $t$:
@@
    the partial route $\pi^{t}=\left(v_{0},v^{1},v^{2},\dots,v^{t-1}\right)$,
@@
We remark that, ideally, a well-trained policy network can automatically extract information about the remaining demand $d^{t}$ given the instance $U$ and the current partial route $\pi^{t}$, but we find that explicitly providing $d^{t}$ to the policy network can effectively reduce the computational load for decision-making and improve the solution quality. Therefore, we include the remaining demand $d^{t}$ as part of the state $s_{t}$.
@@ 3.2.2 Action
At each decision step $t$, an action $a_{t}\in A_{t}$ is defined as selecting the next node (i.e., a new market or the depot) to visit. If a new market is selected, the purchaser will visit it and start the next step, $t+1$. Otherwise, if the depot is selected to visit again, it means that a complete route is formed, and the route construction terminates.
Note that an arbitrarily selected action may lead to infeasible solutions, either because 1) the constructed route may not be a simple cycle, or 2) the markets on the route cannot support a feasible purchasing plan that fulfills the product demands in the second stage. Therefore, to avoid infeasibility, we introduce two masking rules over the action set $A_{t}$ at each decision step $t$. First, to ensure a simple cycle, markets already in the partial route $\pi^{t}$ are excluded from $A_{t}$. In fact, it makes no sense to revisit a market from the perspective of optimization objective, since the traveling cost will increase strictly, but no savings on the purchasing cost can be made. Second, to guarantee the existence of a feasible purchasing plan, we mask the depot from $A_{t}$ until the remaining demand $d_{k}^{t}=0$ for all $k\in K$. In other words, the purchaser is forbidden to return to the depot unless all product demands can possibly be satisfied from the current partial route. These two masking rules ensure that the MDP always produces a feasible route and purchasing plan regardless of the policy.